OpenStudy (anonymous):

How many atoms are in a sample of 175 grams of sodium (Na)?

OpenStudy (abb0t):

\(175~grams~Na^+ \times \frac{1~mol~Na^+}{molar~mass~Na^+}\) now, divide by avogadro's number, which is \(6.022 \times 10^{23} \large \frac{atoms}{mol}\)

OpenStudy (unklerhaukus):

the molar mass of sodium is about 23 grams